what will be the concentration if 0.7 g NaCl is poured in 0.3 liter destilled water?


1
Expert's answer
2021-12-21T12:35:04-0500

CM = n / V

n = m / M

M (NaCl) = 58.44 g/mol

n (NaCl) = 0.7 / 58.44 = 0.012 mol

CM (NaCl) = 0.012 / 0.3 = 0.04 M